Consider the hypothesis test H0: μ1 = μ2 against H1: μ1 > μ2 with known variances σ1=10 and σ2=5. Suppose that sample sizes n1=10 and n2=15 and that x-bar1 = 24.5 and x-bar2 = 21.3. Use alpha = .01. Determine the confidence interval.

a) =0

b) ≥2.78

c) ≥3.04

d) ≥-4.74
Answer #1

Solution:-

The 99% confidence interval is C.I = (- 6.398, 12.798).

C.I = \left ( x_{1} -x_{2}\right )\pm t_{\alpha /2}\times \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}

C.I = 3.2 + 2.81 × 3.41565

C.I = 3.2 + 9.598

C.I = (- 6.398, 12.798)